OpenStudy (lolacole12):

What is the sum? –40 + 23

OpenStudy (lolacole12):

ill give medal

OpenStudy (anonymous):

-17

OpenStudy (lolacole12):

it is

pooja195 (pooja195):

Explain why.

OpenStudy (lolacole12):

please do

OpenStudy (anonymous):

ok so start at -40 add 23 which is -17

OpenStudy (lolacole12):

ohh wich is 17

pooja195 (pooja195):

Good :-)

OpenStudy (anonymous):

dont forget the -

OpenStudy (lolacole12):

-17 thanks i got 1 more

OpenStudy (anonymous):

ok bring it on

OpenStudy (lolacole12):

What is the value of u + v + w when u = –10, v = 4, and w = 8? A. –22 B. 2 C. 14 D. 22

OpenStudy (lolacole12):

and

OpenStudy (lolacole12):

Which statements are true? Choose all answers that are correct. A. –15 + (–10) > –4 + (–20) B. –10 + (–11) > 3 + (–26) C. 17 + (–26) > 12 + (–19) D. –18 + 31 = 24 + (–11)

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Add -10 and 4 to get -6 Then -6+8 B. 2

OpenStudy (anonymous):

hey explain @Muslimah12

OpenStudy (anonymous):

you need to explain it with your answers

OpenStudy (lolacole12):

shes fine shes a great help so are u

OpenStudy (anonymous):

aw thats what i was gonna explain but diffrent

OpenStudy (anonymous):

READ THAT

OpenStudy (anonymous):

i know @Austin1617

OpenStudy (lolacole12):

help the second one this time explain thanks

OpenStudy (anonymous):

then explaination first then answer

OpenStudy (anonymous):

But some question can't be really explained thought

pooja195 (pooja195):

Hi. Let's avoid being rude to each other. Always explain your answers @Muslimah12 :)

OpenStudy (anonymous):

Hey its pooja

pooja195 (pooja195):

Giving answers does not teach the answer anything :[

OpenStudy (anonymous):

Woo Woo

OpenStudy (lolacole12):

Which statements are true? Choose all answers that are correct. A. –15 + (–10) > –4 + (–20) B. –10 + (–11) > 3 + (–26) C. 17 + (–26) > 12 + (–19) D. –18 + 31 = 24 + (–11) ill try to do it to

OpenStudy (anonymous):

Pooja with the kitty image

OpenStudy (lolacole12):

can u help please pooja

OpenStudy (anonymous):

why will no one help me with my math question

OpenStudy (anonymous):

D:

OpenStudy (lolacole12):

b and d

OpenStudy (anonymous):

b and d what

OpenStudy (lolacole12):

Which statements are true? Choose all answers that are correct. A. –15 + (–10) > –4 + (–20) B. –10 + (–11) > 3 + (–26) C. 17 + (–26) > 12 + (–19) D. –18 + 31 = 24 + (–11)

OpenStudy (lolacole12):

i get it